# Vevo2: Bridging Controllable Speech and Singing Voice Generation via Unified Prosody Learning

### v1

@@ 4.3 Controllability of Vevo2
Timbre and Style Control To evaluate Vevo2’s controllability of timbre and style, we conduct various tasks including zero-shot voice conversion (VC), singing voice conversion (SVC), speech style conversion (for accent and emotion), and singing style conversion. The results of zero-shot SVC task are presented in Table 3, with other experimental results provided in Appendix C.
@@
For zero-shot SVC, we include five baselines: three models pre-trained only on speech data (FACodec Ju et al. (2024), Vevo Zhang et al. (2025a), and CosyVoice 2 Du et al. (2024b)), and two models pre-trained on both speech and singing voice data (NeuCoSVC 2 Sha et al. (2024) and SeedVC Liu (2024)). Notably, for Vevo and CosyVoice 2, we utilize their flow-matching (FM) models. Similarly, we include Vevo2’s FM model as a baseline (denoted as Vevo2-FM). The results in Table 3 reveal two key findings: (1) Comparing Vevo2-FM with the existing baselines, Vevo2-FM demonstrates superior timbre-disentanglement (higher SIM for both English and Chinese), validating its effectiveness for timbre control. Additionally, it exhibits strong melody modeling capability (high Melody-MOS), attributable to the explicit modeling of chromagram features in our content-style tokenizer. However, due to the lower frame rate of content-style tokens (12.5 Hz), Vevo2-FM shows slightly lower intelligibility compared to Vevo-FM (50 Hz), CosyVoice2-FM (25 Hz), and SeedVC (using continuous features). (2) Comparing Vevo2-FM and Vevo2, there are two interesting observations: a) Vevo2 shows significant improvements in intelligibility, which we attribute to both the explicit use of text input and intelligibility alignment during post-training. b) Although Vevo2’s Melody-MOS is lower than Vevo2-FM’s, its absolute value still exceeds 2, indicating its ability to “roughly follow the melody contour”. This trade-off benefits style-converted SVC [^3] For style-converted SVC, we aim for the generated singing voice to follow the source’s coarse-grained melody contour while imitating fine-grained expression details (such as vibrato) from the reference Zhang et al. (2025a); Violeta et al. (2025).: specifically, Vevo2 achieves better style imitation performance (higher Style-CMOS), leading to improved speaker similarity (higher SIM). Similar characteristics are also observed in style-converted VC tasks Zhang et al. (2025a).
4.4 Effectiveness of Multi-Objective Post-Training
To explore the individual effects of intelligibility alignment and prosody similarity alignment, we conduct an ablation study on Vevo2’s post-training process. Specifically, based on the pre-trained model (i.e., Vevo2-base), we introduce a post-trained model that exclusively uses intelligibility reward to optimize, denoted as Base w/ Intell. We use two challenging tasks to evaluate: humming-to-singing and instrument-to-singing, which also demonstrate Vevo2’s unique capabilities in prosody control. For the instrument-to-sing task, we design two subjective metrics (see Appendix D for details): given a target text, a target melody (i.e., the instrument sound), and a generated voice, subjects perform binary classification to determine whether the voice accurately sings the text (text accuracy) and the melody (melody accuracy). The results shown in Figure 3 reveal several key findings. When using only the intelligibility reward, the model demonstrates significant improvement in intelligibility (WER and Text Accuracy). However, this single-objective optimization potentially compromises other objectives, as evidenced by Base w/ Intell’s Melody Accuracy dropping even below that of the pre-trained model (from 65.0% to 50.0%). Interestingly, when employing both rewards jointly, we observe not only the enhanced melody-following capability but also the improved text-following ability (e.g., text accuracy increases from 75.0% to 90.0%). We hypothesize that this improvement stems from strengthened prosody modeling, which in turn benefits pronunciation and consequently enhances intelligibility. This further verifies the advantages of our proposed multi-objective post-training approach.
@@ 4.4 Effectiveness of Multi-Objective Post-Training
5 Conclusion
Appendix A Details of Vevo2
A.1 Prosody Tokenizer and Content-Style Tokenizer
This paper processes and trains on speech and singing voice data with a sampling rate of 24 kHz. For chromagram feature extraction, we employ Librosa’s implementation [^4] https://librosa.org/doc/main/generated/librosa.feature.chroma_stft.html with the following parameters: number of FFT points set to 1920, hop length of 480, window size of 1920, and 24 chroma bins. For Whisper hidden features Radford et al. [2023], we utilize the encoder output from the pre-trained whisper-medium model [^5] https://huggingface.co/openai/whisper-medium. We adopt the VQ-VAE implementations from Amphion [^6] https://github.com/open-mmlab/Amphion/tree/main/models/vc/vevo Zhang et al. [2024a], Li et al. [2025], which serves for MaskGCT Wang et al. [2025a] and Vevo Zhang et al. [2025a]. Our only modification is the addition of a CNN-based downsampling layer before the encoder input and a CNN-based upsampling layer after the decoder output. The extracted chromagram features and Whisper features are both at 50 Hz. For the prosody tokenizer, the down(up)-sampling ratio is 8, resulting in a token rate of 6.25 Hz, with a codebook vocabulary size of 512. For the content-style tokenizer, the down(up)-sampling ratio is 4, yielding a token rate of 12.5 Hz, with a codebook vocabulary size of 16,384.
All models in this study use AdamW Kingma and Ba [2015], Loshchilov and Hutter [2019] as the optimizer. For the two tokenizers, we use a peak learning rate of 1e-4, with linear warm-up for 10K steps followed by decay over the remaining training period. Both tokenizers are trained for 300K updates. The prosody tokenizer and content-style tokenizer contain 38M and 44M parameters, respectively.
A.2 Content-Style Modeling
During the content-style modeling stage, we initialize the AR model using Qwen2.5-0.5B [^7] https://huggingface.co/Qwen/Qwen2.5-0.5B Yang et al. [2024b] and expand its original vocabulary to accommodate our proposed prosody tokens and content-style tokens. The total parameter count of the AR model is 509M. The text instruction prefixes for IPL ($\bm{I}_{ipl}$) and EPL ($\bm{I}_{epl}$) are defined as: "User will provide you with a text. Please vocalize it with natural expression." and "User will provide you with a text. Please first generate a good prosodic instruction, then vocalize the text based on it." respectively. For the BPE-based text tokenizer, we adopted the same tokenization strategy as CosyVoice 2, which masks out one-to-many tokens to prevent excessive pronunciation length and mitigate corner cases caused by data sparsity Du et al. [2024b]. During pre-training, we use a peak learning rate of 5e-4, with linear warm-up for 32K steps followed by decay over the remaining training period. The model is pre-trained for 500K updates.
A.3 Acoustic Modeling
During the acoustic modeling stage, we adopted the flow-matching (FM) transformer architecture from Vevo. The transformer consists of 16 layers with 16 attention heads and a hidden dimension of 1024. The FM model contains 363M parameters in total. To enhance training efficiency, we incorporated the REPA strategy Yu et al. [2025] and enforced alignment between the hidden output of the transformer’s 5th layer and the W2v-BERT 2.0 features [^8] https://huggingface.co/facebook/w2v-bert-2.0 Chung et al. [2021], Wang et al. [2025a] during the training process. We use a peak learning rate of 7.5e-5, with linear warm-up for 32K steps followed by decay over the remaining training period. The FM model is trained for 700K updates.
To develop a unified vocoder for both speech and singing voice, we fine-tuned Vevo’s speech vocoder [^9] https://huggingface.co/amphion/Vevo Zhang et al. [2025a] using our training data (comprising 101K hours of speech and 7K hours of singing voice data). The vocoder, which is based on the Vocos architecture Siuzdak [2024], contains 255M parameters and underwent fine-tuning for 572K updates.
A.4 Post-training
(a) Pre-training with EPL and IPL.
(b) Reward model training with EPL and IPL.
Figure 4: Explicit Prosody Learning (EPL) and Implicit Prosody Learning (IPL).
A.4.1 Intelligibility Reward Model
For training an intelligibility reward model, we employed the widely-adopted Bradley-Terry (BT) approach Bradley and Terry [1952], Ouyang et al. [2022], Bai et al. [2022]. As illustrated in Figure 4, we utilized a preference corpus (INTP Zhang et al. [2025b] in our study) to train the pairwise reward model using positive and negative samples (Figure 4(b)). Analogous to our pre-training strategy that incorporated both EPL and IPL strategies (Figure 4(a)), we randomly alternated between IPL and EPL with equal probability during reward model training.
The reward model was initialized using our pre-trained AR model. We implemented a learning rate schedule with a peak value of 1e-4, incorporating a linear warm-up phase for the first 5K steps, followed by a decay phase throughout the remaining training period. The training process continued for 55K updates. During inference, the model accepts either a $\bm{T},\bm{Q}_{cs}]$ or $\bm{T},\bm{Q}{p},\bm{Q}_{cs}]$ sequence as input and produces a float value representing the intelligibility reward.
A.4.2 Prosody Similarity Reward Model
For the prosody similarity alignment, we prepared prompt data incorporating both text and prosody tokens. Specifically, we employed three types of prompts: (1) For speech data, we utilized 20K positive samples randomly selected from INTP, retaining their original text and extracting prosody tokens from their corresponding waveforms. (2) For singing voice data, we selected 20K samples from M4Singer Zhang et al. [2022a], similarly preserving their original text and extracting prosody tokens from the original waveforms. (3) For instrumental sounds, we employed 16 MIDI instruments (Table 4).
Table 4: The 16 MIDI instruments used for post-training.
Category Instrument Name
Piano Acoustic Grand Piano
Bright Acoustic Piano
Electric Piano
Strings Violin
Cello
Orchestral Harp
Woodwinds Alto Sax
Flute
Clarinet
Brass Trumpet
Trombone
Folk Pan Flute
Koto
Voice Choir Aahs
Voice Oohs
Synth Voice
During each training iteration with the 20K M4Singer samples, we randomly selected one instrument and rendered the corresponding audio using pretty_midi [^10] https://github.com/craffel/pretty-midi. We use the original text of the sample and extract prosody tokens from these rendered instrumental sounds.
A.5 Inference-Time Control
A.5.1 Duration Control
@@
As discussed in Section 3.5, we discovered that Vevo2 enables effective duration control of generated outputs through the utilization of prosody tokens as input. This capability arises from a fundamental characteristic: both our proposed prosody tokenizer and content-style tokenizer operate at fixed frame rates (6.25 Hz and 12.5 Hz, respectively). During training, this results in a consistent linear relationship between the sequence lengths of content-style tokens and prosody tokens, specifically maintaining a 2:1 ratio. The AR model readily learns this straightforward length pattern. Consequently, during inference, we can control the output duration by applying linear scaling transformations to the chromagram features of the prosodic source—leveraging the inherent robustness of chromagram features to such transformations—thereby modulating the length of prosody tokens.
Table 5: Effectiveness of duration control of Vevo2 for different tasks.
Task DDUR ($\downarrow$) Duration Consistency ($\uparrow$)
Speech Editing 0.22s 97.1%
Singing Lyric Editing 0.20s 98.4%
Singing Voice Conversion (style-converted) 0.12s 98.4%
@@
In our experiments, we investigated Vevo2’s duration control effectiveness across three tasks: speech editing, singing lyric editing, and singing voice conversion. We adopt the DDUR metric Zhang et al. [2020, 2025a] ($\downarrow$), which measures the average absolute difference in seconds between the actual generated duration ($\hat{d}$) and the target duration ($d$), defined as $\text{abs}(\hat{d}-{d})$. Additionally, we introduce duration consistency ($\uparrow$), defined as $\frac{\text{abs}(\hat{d}-{d})}{d}$, to evaluate the relative accuracy of duration control. For editing tasks, following F5-TTS Chen et al. [2025], we calculate the target duration based on the linear relationship between the lengths of edited and raw text. For singing voice conversion, we constrain the target duration to match the source waveform duration. As demonstrated in Table 5, Vevo2 achieved over 97% duration consistency across all three tasks, exhibiting good duration controllability despite its autoregressive nature.
A.5.2 Pitch Region Control
@@
The pitch region control is particularly effective for voice conversion (VC) and singing voice conversion (SVC) tasks Zhang et al. [2024e], Ning et al. [2023], Huang et al. [2023]. For Vevo2, to achieve the pitch region control:
@@
Table 6: Effect of pitch shift of Vevo2 in VC and SVC tasks.
@@
The effectiveness of Vevo2’s pitch shift control in VC and SVC tasks is demonstrated in Table 6. The results reveal two key findings: First, using pitch shift significantly enhances speaker similarity (SIM) for both Vevo2-FM (using only the FM model) and the complete Vevo2 system (utilizing both AR and FM models). Second, we observe a slight increase in WER when applying pitch shift. We hypothesize that this minor degradation occurs because the content-style tokens and prosody tokens extracted from inference-time pitch-shifted audio represent patterns less frequently encountered in the training data. In future work, we plan to incorporate such pitch shift perturbations during training to mitigate these small discrepancies arising from the train-inference mismatch.
Appendix B Details of Baselines
@@
  • F5-TTS Chen et al. [2025]: It adopts a single-stage flow-matching transformer to predict Mel spectrograms from text. It is pre-trained on Emilia He et al. [2024] (101K hours of speech data). We use the officially released checkpoint [^11] https://huggingface.co/SWivid/F5-TTS to generate samples.
  • MaskGCT Wang et al. [2025a]: It adopts a two-stage pipeline: text to content-style tokens (50 Hz), and content-style tokens to acoustic tokens. It is pre-trained on Emilia He et al. [2024] (101K hours of speech data). We use the officially released checkpoint [^12] https://huggingface.co/amphion/MaskGCT to generate samples.
  • CosyVoice 2 Du et al. [2024a]: It contains an AR transformer to predict the content-style tokens (25 Hz) from text, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^13] https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B to generate samples, which is pre-trained on 167K hours of private data.
@@
  • StyleSinger Zhang et al. [2024d]: It is a zero-shot SVS model for style transfer that is pre-trained on the Chinese GTSinger data Zhang et al. [2024b]. We use the officially released checkpoint [^14] https://github.com/AaronZ345/StyleSinger to generate samples.
  • TCSinger Zhang et al. [2024c]: It is a zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. It is pre-trained on the GTSinger data Zhang et al. [2024b]. We use the officially released checkpoint [^15] https://github.com/AaronZ345/TCSinger to generate samples.
@@
Zero-Shot Voice Conversion and Singing Voice Conversion
  • FACodec Ju et al. [2024]: It adopts an auto-encoder and residual vector quantization based architecture. It decouples the raw waveform into factorized attributes through ASR, F0 prediction, and speaker classification tasks, trained on the Libri-light dataset Kahn et al. [2020]. We use the released checkpoint in Amphion  [^19] https://huggingface.co/amphion/naturalspeech3_facodec Zhang et al. [2024a], Li et al. [2025] (which is implemented by the authors) to generate samples.
  • Vevo-FM Zhang et al. [2025a]: It is the flow-matching model of Vevo (i.e., Vevo-Timbre in the original paper), which uses a 50 Hz VQ-VAE content-style tokenizer based on HuBERT Hsu et al. [2021] representations. It is pre-trained on Emilia (101K hours of speech data). We use the officially released checkpoint [^20] https://huggingface.co/amphion/Vevo to generate samples.
  • CosyVoice2-FM Du et al. [2024b]: It is the flow-matching model of CosyVoice 2 that uses 25 Hz content-style tokenizer, consistent with the configuration described in the zero-shot TTS task above.
  • NeuCoSVC 2 Sha et al. [2024]: It is an enhanced version of NeuCoSVC and is pre-trained on an extensive internal dataset comprising approximately 500 hours of singing voice data and various open-source speech datasets Sha et al. [2024]. We use the officially released checkpoint [^21] https://github.com/thuhcsi/NeuCoSVC/tree/NeuCoSVC2 to generate samples.
  • SeedVC Liu [2024]: It is a diffusion transformer that uses an external timbre shifter during training to perturb the source speech timbre. We use the officially released pre-trained VC and SVC models [^22] https://github.com/Plachtaa/seed-vc to conduct the inference.
Zero-Shot Emotion Conversion and Accent Conversion
  • Vevo Zhang et al. [2025a]: It is named as Vevo-Style in the original paper, which contains an AR transformer to predict content-style tokens (50 Hz) from content tokens, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^23] https://huggingface.co/amphion/Vevo to generate samples, which is pre-trained on 101K hours of Emilia.
Appendix C Additional Experimental Details and Results
C.1 Zero-Shot Text to Speech
The zero-shot TTS evaluation results for expressive speech and singing voice domains are presented in Table 1. Table 7 demonstrates the performance on the regular speech domain using the SeedTTS evaluation sets Anastassiou et al. [2024a], a widely adopted benchmark in the TTS field Wang et al. [2025a], Du et al. [2024b], Chen et al. [2025]. The results corroborate our findings from Section 4.2. Specifically, we observe that incorporating singing voice data alongside speech data enhances performance across all metrics (WER, SIM, UTMOS) in the regular speech domain. Notably, the post-trained Vevo2 achieves comparable performance to existing state-of-the-art baselines, despite utilizing a content-style tokenizer with a lower frame rate.
Table 7: Results of zero-shot TTS task on regular speech domain. The results on expressive speech and singing voice domains are shown in Table 1. (PT?: Whether post-training is applied.)
@@ C.1 Zero-Shot Text to Speech
F5-TTS Chen et al. [2025] - 101K 3.442 0.670 3.090
MaskGCT Wang et al. [2025a] 50 101K 2.344 0.738 3.136
CosyVoice2 Du et al. [2024a] 25 167K 2.091 0.709 3.685
@@
C.2 Zero-Shot Editing
The zero-shot editing evaluation (Table 2) was conducted using evaluation sets constructed with DeepSeek-V3 DeepSeek-AI et al. [2024]. Specifically, we utilized the same expressive speech and singing voice evaluation data from our zero-shot TTS tasks. For each sample’s text, we generated edited versions using DeepSeek-V3 with the following prompt: "Please edit the following text by adding, deleting, or modifying content as needed. Note: (1) Ignore any existing punctuation errors in the original text and ensure proper punctuation in the edited version. (2) The edited version must differ from the original text. (3) Return only the final edited text without any additional explanations."
C.3 Zero-Shot Voice Conversion
Table 8 presents the results of our voice conversion evaluation, utilizing the same baselines [^24] For SeedVC, we employed the official SVC checkpoint for the zero-shot SVC task (Table 3) and the official VC checkpoint for the zero-shot VC task (Table 8). as the SVC task in Table 3, where we use the SeedTTS VC evaluation set Anastassiou et al. [2024a]. The VC task results lead to conclusions similar to those observed in the SVC experiments: First, Vevo2-FM demonstrates robust timbre disentanglement capabilities, with speaker similarity (SIM) metrics matching or exceeding existing VC models, particularly for English content. Furthermore, the transition from Vevo2-FM to Vevo2 (i.e., from style-preserved VC to style-converted VC) yields additional improvements in speaker similarity. Moreover, the explicit incorporation of text input significantly enhances intelligibility performance.
Table 8: Results on zero-shot VC task.

### v2

@@ V-C Controllability over Timbre and Style
To evaluate Vevo2’s controllability of timbre and style, we conduct tasks including zero-shot VC (see the supplementary materials), SVC, and speech style conversion (for accent and emotion).
Zero-Shot SVC For zero-shot SVC, we include five baselines: three models pre-trained only on speech data (FACodec [naturalspeech3], Vevo [vevo], and CosyVoice 2 [cosyvoice2]), and two models pre-trained on both speech and singing voice data (NeuCoSVC 2 [neucosvc] and SeedVC (SVC) [seedvc]). Notably, for Vevo and CosyVoice 2, we utilize their flow-matching (FM) models. Similarly, we include Vevo2’s FM model as a baseline (denoted as Vevo2-FM). The results in Table III reveal two key findings:
(1) Comparing Vevo2-FM with the existing baselines, Vevo2-FM demonstrates superior timbre-disentanglement (higher SIM for both English and Chinese), validating its effectiveness for timbre control. Additionally, it exhibits strong melody modeling capability (high Melody-MOS), attributable to the explicit modeling of chromagram features in our content-style tokenizer. However, due to the lower frame rate of content-style tokens (12.5 Hz), Vevo2-FM shows slightly lower intelligibility compared to Vevo-FM (50 Hz), CosyVoice2-FM (25 Hz), and SeedVC (using continuous features).
@@
(2) Comparing Vevo2-FM and Vevo2, there are two interesting observations: a) Vevo2 shows significant improvements in intelligibility, which we attribute to both the explicit use of text input and intelligibility alignment during post-training. b) Although Vevo2’s Melody-MOS is lower than Vevo2-FM’s, its absolute value still exceeds 2, indicating its ability to “roughly follow the melody contour”. This trade-off benefits style-converted SVC [^12] For style-converted SVC, we aim for the generated singing voice to follow the source’s coarse-grained melody contour while imitating fine-grained expression details (such as vibrato) from the reference [vevo, svcc2025].: specifically, Vevo2 achieves better style imitation performance (higher Style-CMOS), leading to improved speaker similarity (higher SIM). Similar characteristics are also observed in style-converted VC tasks [vevo].
Zero-Shot Accent and Emotion Conversion Table V demonstrates Vevo2’s performance on zero-shot accent and emotion conversion tasks. We benchmarked against Vevo [vevo] and adopted their evaluation sets and two style conversion metrics: Accent-SIM (computed using a pre-trained accent classifier [^13] https://huggingface.co/Jzuluaga/accent-id-commonaccent_ecapa) and Emotion-SIM (calculated via a pre-trained emotion classifier [^14] https://github.com/ddlBoJack/emotion2vec) to assess style conversion performance. The results indicate that Vevo2 achieves comparable style similarity scores to Vevo while exhibiting superior performance in both intelligibility and speaker similarity metrics.
@@
V-D Controllability over Duration
@@
As discussed in Section III-D, we discovered that Vevo2 enables effective duration control of generated outputs through the utilization of prosody tokens as input. This capability arises from a fundamental characteristic: both our proposed prosody tokenizer and content-style tokenizer operate at fixed frame rates (6.25 Hz and 12.5 Hz, respectively). During training, this results in a consistent linear relationship between the sequence lengths of content-style tokens and prosody tokens, specifically maintaining a 2:1 ratio. The AR model readily learns this straightforward length pattern. Consequently, during inference, we can control the output duration by applying linear scaling transformations to the chromagram features of the prosodic source—leveraging the inherent robustness of chromagram features to such transformations—thereby modulating the length of prosody tokens.
@@
In our experiments, we investigated Vevo2’s duration control effectiveness across three tasks: speech editing (SE), singing lyric editing (SLE), and style-converted SVC. We adopt the metric, Differences of Duration (DDUR, $\downarrow$) [non-parallel-seq2seq-vc, vevo], which measures the average absolute difference in seconds between the actual generated duration ($\hat{d}$) and the target duration ($d$). Additionally, we introduce Duration Consistency (DC, $\uparrow$) to evaluate the relative accuracy of duration control:
$\textbf{DDUR}=\text{abs}(\hat{d}-{d}),\quad\textbf{DC}=\frac{\text{abs}(\hat{d}-{d})}{d}.$ (7)
For editing tasks, following F5-TTS [f5tts], we calculate the target duration based on the linear relationship between the lengths of edited and raw text. For the style-converted SVC, we constrain the target duration to match the source waveform duration. As demonstrated in Table V, Vevo2 achieved over 97% duration consistency across all three tasks, exhibiting good duration controllability despite its autoregressive nature.
V-E Controllability over Pitch Region
@@
The pitch region control is particularly effective for voice conversion (VC) and singing voice conversion (SVC) tasks [amphion-svc, svcc-vits-ziqian, svcc-2023]. For Vevo2, to achieve the pitch region control:
@@
TABLE VI: Effect of pitch region control of Vevo2.
@@
The effectiveness of Vevo2’s pitch shift control in VC and SVC tasks is demonstrated in Table VI. The results reveal two key findings: First, using pitch shift significantly enhances speaker similarity (SIM) for both Vevo2-FM (using only the FM model) and the complete Vevo2 system (utilizing both AR and FM models). Second, we observe a slight increase in WER when applying pitch shift. We hypothesize that this minor degradation occurs because the content-style tokens and prosody tokens extracted from inference-time pitch-shifted audio represent patterns less frequently encountered in the training data. In future work, we plan to incorporate such pitch shift perturbations during training to mitigate these small discrepancies arising from the train-inference mismatch.
V-F Effectiveness of Multi-Objective Post-Training
To explore the individual effects of intelligibility alignment and prosody similarity alignment, we conduct an ablation study on Vevo2’s post-training process. Specifically, based on the pre-trained model (i.e., Vevo2-base), we introduce a post-trained model that exclusively uses intelligibility reward to optimize, denoted as Base w/ Intell. We use two challenging tasks to evaluate: humming-to-singing and instrument-to-singing, which also demonstrate Vevo2’s unique capabilities in prosody control. For the instrument-to-sing task, we design two subjective metrics (see the supplementary materials for details): given a target text, a target melody (i.e., the instrument sound), and a generated voice, subjects perform binary classification to determine whether the voice accurately sings the text (text accuracy) and the melody (melody accuracy).
The evaluation sets for humming-to-singing and instrument-to-singing tasks are constructed through the following procedure. For melody sources, we selected humming samples [humtrans] and instrumental sounds, including flute [flute-dataset] and violin [violin-dataset] recordings. For each melody sample, we leveraged its MIDI annotation’s note count to generate appropriate lyrics using DeepSeek-v3. The following prompt was used for DeepSeek-v3: "I will provide the number of notes in a melody. Please write suitable lyrics for it. Note: (1) The number of words in the lyrics should roughly match the number of notes; (2) Please only return the result, without any further explanation." After obtaining both melody and lyrics, we randomly assigned a target singer from SingStyle111 [singstyle111] for each sample, thereby creating evaluation triplets in the form of (melody, lyric, singer). The evaluation set comprises 1,000 samples in total, evenly distributed between English (500) and Chinese (500) lyrics.
The results shown in Figure 5 reveal several key findings. When using only the intelligibility reward, the model demonstrates significant improvement in intelligibility (WER and Text Accuracy). However, this single-objective optimization potentially compromises other objectives, as evidenced by Base w/ Intell’s Melody Accuracy dropping even below that of the pre-trained model (from 65.0% to 50.0%). Interestingly, when employing both rewards jointly, we observe both the enhanced melody-following and text-following ability (e.g., text accuracy increases from 75.0% to 90.0%). We hypothesize that this improvement stems from strengthened prosody modeling, which in turn benefits pronunciation and consequently enhances intelligibility. This further verifies the advantages of our proposed multi-objective post-training.
VI Conclusion
Appendix A Details of Baselines
@@
  • F5-TTS [f5tts]: It adopts a single-stage flow-matching transformer to predict Mel spectrograms from text. It is pre-trained on Emilia [emilia] (101K hours of speech data). We use the officially released checkpoint [^15] https://huggingface.co/SWivid/F5-TTS to generate samples.
  • MaskGCT [maskgct]: It adopts a two-stage pipeline: text to content-style tokens (50 Hz), and content-style tokens to acoustic tokens. It is pre-trained on Emilia [emilia] (101K hours of speech data). We use the officially released checkpoint [^16] https://huggingface.co/amphion/MaskGCT to generate samples.
  • CosyVoice 2 [cosyvoice]: It contains an AR transformer to predict the content-style tokens (25 Hz) from text, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^17] https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B to generate samples, which is pre-trained on 167K hours of private data.
@@
  • StyleSinger [stylesinger]: It is a zero-shot SVS model for style transfer that is pre-trained on the Chinese GTSinger data [gtsinger]. We use the officially released checkpoint [^18] https://github.com/AaronZ345/StyleSinger to generate samples.
  • TCSinger [tcsinger]: It is a zero-shot SVS model for style transfer across cross-lingual speech and singing styles, along with multi-level style control. It is pre-trained on the GTSinger data [gtsinger]. We use the officially released checkpoint [^19] https://github.com/AaronZ345/TCSinger to generate samples.
@@
TABLE VII: Results of zero-shot TTS task on regular speech domain. The results on expressive speech and singing voice domains are shown in Table I. (PT?: Whether post-training is applied.)
@@
F5-TTS [f5tts] - 101K 3.442 0.670 3.090
MaskGCT [maskgct] 50 101K 2.344 0.738 3.136
CosyVoice2 [cosyvoice] 25 167K 2.091 0.709 3.685
@@
Zero-Shot Voice Conversion and Singing Voice Conversion
  • FACodec [naturalspeech3]: It adopts an auto-encoder and residual vector quantization based architecture. It decouples the raw waveform into factorized attributes through ASR, F0 prediction, and speaker classification tasks, trained on the Libri-light dataset [libri-light]. We use the released checkpoint in Amphion  [^23] https://huggingface.co/amphion/naturalspeech3_facodec [amphion, amphion_v0.2] (which is implemented by the authors) to generate samples.
  • Vevo-FM [vevo]: It is the flow-matching model of Vevo (i.e., Vevo-Timbre in the original paper), which uses a 50 Hz VQ-VAE content-style tokenizer based on HuBERT [hubert] representations. It is pre-trained on Emilia (101K hours of speech data). We use the officially released checkpoint [^24] https://huggingface.co/amphion/Vevo to generate samples.
  • CosyVoice2-FM [cosyvoice2]: It is the flow-matching model of CosyVoice 2 that uses 25 Hz content-style tokenizer, consistent with the configuration described in the zero-shot TTS task above.
  • NeuCoSVC 2 [neucosvc]: It is an enhanced version of NeuCoSVC and is pre-trained on an extensive internal dataset comprising approximately 500 hours of singing voice data and various open-source speech datasets [neucosvc]. We use the officially released checkpoint [^25] https://github.com/thuhcsi/NeuCoSVC/tree/NeuCoSVC2 to generate samples.
  • SeedVC [seedvc]: It is a diffusion transformer that uses an external timbre shifter during training to perturb the source speech timbre. We use the officially released pre-trained VC and SVC models [^26] https://github.com/Plachtaa/seed-vc, which are denoted as SeedVC (VC) and SeedVC (SVC), to conduct the inference.
Zero-Shot Emotion Conversion and Accent Conversion
  • Vevo [vevo]: It is named as Vevo-Style in the original paper, which contains an AR transformer to predict content-style tokens (50 Hz) from content tokens, and a flow-matching transformer to predict Mel spectrograms. We use the officially released checkpoint [^27] https://huggingface.co/amphion/Vevo to generate samples, which is pre-trained on 101K hours of Emilia.
Appendix B Additional Experimental Results
B-A Zero-Shot Text to Speech
The zero-shot TTS evaluation results for expressive speech and singing voice domains are presented in Table I. Table VII demonstrates the performance on the regular speech domain using the SeedTTS evaluation sets [seedtts], a widely adopted benchmark in the TTS field [maskgct, cosyvoice2, f5tts]. The results corroborate our findings from Section V-A. Specifically, we observe that incorporating singing voice data alongside speech data enhances performance across all metrics (WER, SIM, UTMOS) in the regular speech domain. Notably, the post-trained Vevo2 achieves comparable performance to existing state-of-the-art baselines, despite utilizing a content-style tokenizer with a lower frame rate.
B-B Zero-Shot Voice Conversion
Table VIII presents the results of our voice conversion evaluation, utilizing the same baselines [^28] For SeedVC, we employed the official SVC checkpoint for the zero-shot SVC task (Table III) and the official VC checkpoint for the zero-shot VC task (Table VIII). as the SVC task in Table III, where we use the SeedTTS VC evaluation set [seedtts]. The VC task results lead to conclusions similar to those observed in the SVC experiments: First, Vevo2-FM demonstrates robust timbre disentanglement capabilities, with speaker similarity (SIM) metrics matching or exceeding existing VC models, particularly for English content. Furthermore, the transition from Vevo2-FM to Vevo2 (i.e., from style-preserved VC to style-converted VC) yields additional improvements in speaker similarity. Moreover, the explicit incorporation of text input significantly enhances intelligibility performance.
TABLE VIII: Results on zero-shot VC task.